Recent Log Entries

HYST AREA 5-2 Cycles Magnets

Log Number: 3626663Logbook(s): Body: 

HYST AREA Version 5-2 has successfully cycled the following:

Area.Linacs.North
Area.Linacs.South
Area.Arcs.Arc1.Spreader
Area.Arcs.Arc1.Arc
Area.Arcs.Arc1.Recombiner
Area.Arcs.Arc1.Extraction
Area.Arcs.Arc2.Spreader
Area.Arcs.Arc2.Arc
Area.Arcs.Arc2.Recombiner
Area.Arcs.Arc2.Extraction
Area.Arcs.Arc3.Spreader
Area.Arcs.Arc3.Arc
Area.Arcs.Arc3.Recombiner
Area.Arcs.Arc3.Extraction
Area.Arcs.Arc4.Spreader
Area.Arcs.Arc4.Arc
Area.Arcs.Arc4.Recombiner
Area.Arcs.Arc4.Extraction

Beam stopper modification

Log Number: 3626662Logbook(s): Body: 

The east arc beam stopper software was modified so that the stopper is not automatically inserted on ioc reboot. This should allow iocea2 to be rebooted more readily in future. The same change will be made for other beam stoppers.

Autolog save

Log Number: 3626658Logbook(s): Tags: Body: saver ran from opsl04

Saved: grames 11-15-18 13:19:59 {A:2 B:5 C:5 E:1050 A=BS B=BS C=CW D=BS 0R08=0.00 0R08 dump in.

iocea2 restored

Log Number: 3626653Logbook(s): Tags: Body: restorelib 1.0 Target restore date was 2018-11-15 13:07:33 The output from the myRestore execution was as follows: iocea2 Execute: /cs/opshome/scripts/60_second_wait.sh (success) Pause: 60 seconds beamStopperEast 2(2) restored blmset 16(16) restored vac 43(43) restored system 3(3) restored Execute: /cs/opshome/scripts/scan_gain_set.sh (success) Pause: 10 seconds bpm 772(772) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocea2 (success)

Hall C slow lock and big BPMs

Log Number: 3626626Logbook(s): Entry Makers: beneschBody: 

is compensating for the horizontal motion caused by fifth pass RF separator. Hall C won't let me turn off the lock (tsk, tsk) so I'll hunt for a time when it was off. Found one: 0100-0215 this morning, figure 3. Top shows the horizontal motion on 3C12 and 3H07C, the last regular BPM. Correctors are at the bottom of the top chart and are fixed. Bottom of figure 3 shows IPM3H07C.XPOS again and the responses of the two big BPMs. The next four charts show 3H07C and each of the four big BPM signals paired. Current was stable at 40 uA CW throughout, figure 8.

Task list for 1pm

Log Number: 3626607Logbook(s): Tags: Entry Makers: michaelmBody: 

-Spot move for decreasing QE (Grames) followed by INJ setup verification(Kazimi)
-IOCEA2 reboot for memory block (Ops); drops in E. Arc beam stopper and so box supplies must be taken down to avoid drop
-5th pass H. Sep. resistor swap (LLRF); requires setup verification by Ops afterwards.
-2L13 CPS blower maintenance (Facilities)
-2L13 2.5 W amp PS swap (EES-RF)
-1L14-1 Tuner frozen; attempting card swap (EES-RF)
-Hall C Kinematics change
-Hall D TAC run (before) / convergence adjustment (after)
-Other tasks? (Call crew chief at x7045 to coordinate)

IHA1C12 Harp Scan w/B Parameter =0.04

Log Number: 3626597Logbook(s): Entry Makers: lrichardBody: 

I originally had the Peak Fitting B Parameter set to 0.35 but then found out that was too high. Advised to increase only in 0.02 increments. Once increased to 0.02 and refitted, fit did not work. Raised B Parameter to 0.04 (Fig. 2) and fit (Fig. 1) was much better than the previous fit.

Operator SIGN IN

Log Number: 3626595Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Moser
DATE:             11/15/2018
SHIFT:               Day

This stamp certifies that the above operator has read/reviewed/received the following:

COMMENT ON UITF Viewers

Log Number: 3626586Logbook(s): Body: 

Cables Pulled

ITVM905- Limit Switch, camera power, solenoid, light, and
video.

IFYM905- Limit Switch and solenoid

Pages